# Learning from Emotions, Demographic Information and Implicit User Feedback in Task-Oriented Document-Grounded Dialogues

### v1

@@ 5 FEDI Analysis
 +User Reaction
43.4 (–0.6) 22.1 (+2.1) 88.2 (–0.1) 96.9 (+10.2) 95.3 (+9.4) 83.5 (+28.7) 77.2 (+16.3) 0.02 60.2 (+7.5)
GPT-2 Feedback-Free GPT-2 34.9 10.4 87.1 88.3 81.6 78.7 69.6 0.02 28.1
 +Emotions 35.1 (+0.2) 10.4 87.1 84.1 (–4.2) 83.8 (+2.2) 75.4 (–3.3) 67.3 (–2.3) 0.02 26.7 (–1.4)
 +Demographics 34.6 (–0.3) 10.4 87.1 80.2 (–0.1) 80.2 (–1.4) 69.3 (–9.4) 57.5 (–12.1) 0.02 26.3 (–1.8)
 +Emotions
 +Demographics
36.0 (+1.1) 11.4 (+1.0) 87.3 (+0.2) 85.1 (–3.2) 84.8 (+3.2) 71.6 (–7.1) 66.7 (– 2.9) 0.02 29.2 (+1.1)
Feedback  +Generation Error 29.2 (–5.7) 8.0 (–2.4) 86.2 (–0.9) 92.4 (+4.1) 91.7 (+10.1) 84.3 (+5.6) 79.3 (–9.7) 0.02 30.9 (+2.8)
 +User Reaction 30.0 (–4.9) 8.3 (–2.1) 86.3 (–0.8) 98.9 (+10.6) 96.5 (+14.9) 83.0 (+4.3) 80.3 (+10.7) 0.02 32.3 (+4.2)
@@
 +User Reaction
30.3 (–4.6) 9.7 (–0.7) 86.4 (–0.7) 94.7 (+6.4) 93.3 (+11.7) 88.0 (+9.3) 80.8 (+11.2 0.01 (–0.01) 35.5 (+7.4)
LLaMA-2 Feedback-Free LLaMA-2 29.3 7.1 86.1 85.9 81.2 37.6 39.2 0.02 28.3
 +Emotions 36.3 (+7.0) 14.9 (+7.8) 85.4 (–0.7) 89.3 (+3.4) 85.3 (+4.1) 40.2 (+2.6) 41.3 (+2.1) 0.01 (–0.01) 18.7 (–9.6)
 +Demographics 33.8 (+4.5) 4.5 (–2.6) 86.5 (+0.4) 85.6 (–0.3) 82.5 (+1.3) 37.1 (–0.5) 40.1 (+0.9) 0.02 21.3 (–7.0)
 +Emotions
 +Demographics
28.8 (–0.5) 5.6 (–1.5) 81.3 (–4.8) 86.7 (+1.8) 87.9 (+6.7) 41.4 (+3.8) 39.6 (+0.4) 0.03 (+0.01) 20.6 (–7.7)
Feedback  +Generation Error 24.1 (–5.2) 7.9 (+0.8) 77.4 (–8.7) 93.1 (+7.2) 95.7 (+14.5) 54.8 (+17.2) 59.6 (+20.4) 0.01 (–0.01) 29.1 (+0.8)
 +User Reaction 24.5 (–4.8) 6.9 (–0.2) 78.8 (–7.3) 94.9 (+9.0) 93.2 (+12.0) 63.5 (+25.9) 70.1 (+30.9) 0.02 27.1 (–1.2)
@@
 +User Reaction
25.0 (–4.3) 9.2 (+1.9) 80.1 (–6.0) 82.4 (–3.5) 83.6 (+2.4) 46.3 (+8.7) 47.2 (+8.0) 0.03 (+0.01) 33.5 (+5.0)
Table 3: Results of our experiments. We use the baseline models as deltas, i.e., the pretrained models finetuned on the FEDI feedback-free dialogues. The models with the greatest improvements are printed in bold.
6 Human Curation Study
To assess the quality of our generated dialogues, we used INCEpTION Klie et al. (2018) as a platform and asked two participants from our test data collection to assess and curate the intent, slot and emotion annotations in 480 feedback-free dialogues and the error and user reaction type annotations in 380 feedback dialogues. To measure the agreement between both annotators (IAA), we use Krippendorff’s Alpha Krippendorff (2006) as provided in the INCEpTION platform.
Annotation Type Missing Changed IAA
Feedback-Free Dialogues Intent 6% 35% 0.90
Slots 56% 19% 0.83
Emotion 2% 81% 0.91
Feedback Dialogues Error Type 16% 36% 0.97
User Reaction Type 16% 34% 0.89
Table 4 shows the results [^12] Overall, 26 dialogues were reported as off-topic and are not considered in these results.. Overall, the ratio of dialogues with at least one missing annotation is low, except for slot annotations. We found that most of them are parcel shipping dialogues, which has a comparatively complex annotation scheme (see Appendix A). We attribute this to our observation on slot and intent annotations from Section 5, i.e., the (occassional) dependence on the background story. A detailed analysis revealed that an average of 1.8 annotations were added to the dialogues, most of them (36%) were slot annotations. For the dialogues with at least one changed annotation, we found that in many of these cases placeholders, e.g., the slot name put in brackets ([shipping_box_name]), were used instead of the slot values from the dialogues (reported by the students). Emotion is the most frequently changed annotation type (on average 2.09 times per affected dialogue), with the originally-annotated emotion often being Neutral.
@@ 7 Experiments and Results
To investigate the impact of combining demographic information with user emotion and implicit feedback, we conduct experiments using three models of different architecture and pretraining approaches, including FLAN-T5 Chung et al. (2022) (780M), GPT-2 Radford et al. (2019) (780M) and LLaMA-2 Touvron et al. (2023) (7B) [^13] The model weights for FLAN-T5 and GPT-2 are available in the Huggingface Model Hub (last accessed 04 January 2024). Access to the weights for LLaMA-2 must be requested from Meta AI (last acessed 26 December 2023).. For LLaMA-2, we only finetune the LoRA Hu et al. (2022) weights in our experiments. Table 3 shows the results achieved in the test data averaged over three runs.
Experimental Setup
We first finetune the pretrained models to our scenario using the feedback-free dialogues (Feedback-Free in Table 3) and included the demographic information (+Demographics) and user emotions (+Emotions) as part of the input sequences. We then use the best performing feedback-free models (the bold ones) for experiments using the feedback dialogues (Feedback). We provide more details, including hyperparameters and input sequences, in Appendix F.
Evaluation Metrics
@@
We use F1-Score (calculated based on overlapping tokens in target and prediction), BLEU(-n) Papineni et al. (2002) and BertScore Zhang et al. (2020) to evaluate the concordance (Matching) of the generated system utterances with the target values. To evaluate for task completion, we use Inform and Success as proposed by Budzianowski et al. (2018) and additionally measure the correctness of the predicted intents (Intent Accuracy) and slot values (Slot Accuracy). To measure the toxicity in the generated responses, we use Perspective API. To measure the factual consistency in the case of question answering, we use $\text{Q}^{2}$ Honovich et al. (2021) [^14] For BLEU and BertScore, we use the implementation from the HuggingFace evaluation library v0.4.1 (last accessed 26 December 2023) and with $n=4$ for BLEU. For Inform and Success, we use the implementation from Nekvinda and Dušek (2021) as a reference (last accessed 28 December 2023). For $Q^{2}$, we use the reference implementation which is available in GitHub (last accessed 29 December 2023). Perspective API is a free-to-use service provided by Google and Jigsaw. Model and training details can be found here (last accessed 04 January 2024)..
Results
In general, we find that including user emotions (+Emotions) has a positive impact. This is particularly significant in the case of LLaMA-2 Touvron et al. (2023), where we used instructions for finetuning that provide additional context, i.e., a brief description of how the user emotion should be considered (see Appendix F for input sequences). In the feedback experiments, we observe great improvements in task completion and factual consistency of generated responses ($Q^{2}$ metric), which are both crucial for task-oriented document-grounded dialogues. For example, by including the generation error in the case of LLaMA-2 Touvron et al. (2023) or by combining the generation error with the user reaction in the case of FLAN-T5 Chung et al. (2022) and GPT-2 Radford et al. (2019). In general, we attribute these improvements to the additional context provided by the generation error and the user reaction, which can be interpreted as a negative example for a response in the specific dialogue context.
@@ Results
We conduct a small-scale human evaluation with two participants from our lab [^15] One of the authors and an intern of our lab who participated during his working hours.. We use the two best feedback-trained models from Table 3 (FLAN-T5 and GPT-2 with generation error and user reaction) and their feedback-free counterparts (FLAN-T5 and GPT-2 with user emotions) to generate responses for 50 randomly chosen samples from the FEDI test set. We asked the two annotators to rate the generated responses for human-likeness (naturalness), relevancy in the dialogue context (coherence), social acceptability (safety), factual consistency (with the target document in the case of answering questions), and engagement (whether they would use this model in practice). We use a likert scale from 1 (lowest rating) to 5 (highest rating) for each attribute and provide the annotators with the knowledge document, dialogue context, and generated response for this evaluation. Table 5 shows the results.
Experiment Naturalness Coherence Safety Engagement
Factual
Consistency
@@
Feedback-Free
4.14 4.15 4.55 3.75 2.20
Feedback
4.30 (+0.16) 4.25 (+0.10) 4.59 (+0.09) 3.89 (+0.14) 2.24 (+0.04)
@@
Feedback-Free
4.24 3.82 4.45 3.44 1.55
Feedback
4.42 (+0.18) 4.05 (+0.23) 4.46 (+0.01) 3.76 (+0.32) 1.58 (+0.03)
Table 5: Results of our human evaluation.
The reponses generated by the feedback-trained FLAN-T5 model are mostly rated higher by the participants. They also reported that the generated responses encourage for more user interaction, e.g., by requesting additional information or paying more attention to the user and their situation, and are in general more factual consistent. In the case of question answering, the generated responses are mostly summaries of the respective documents. According to the authors, the GPT-2 model trained only with emotions already produced very engaging answers, although they are not as coherent as the responses generated by FLAN-T5. This also affects factual consistency in the case of question answering, which is much lower for both GPT-2 models than for FLAN-T5.
@@ 8 Conclusion
In this work, we introduced FEDI, the first dataset for learning from demographic information, user emotions and implicit feedback in task-oriented document-grounded dialogues. FEDI consists of 8.8k dialogues generated and annotated using GPT-3.5-Turbo, 6k of which with annotations for implicit user feedback. For testing, we additionally provide a set of dialogues collected by humans. Our analysis shows the effectiveness of our generation framework and that FEDI is comparable to other related datasets in terms of size, utterance length, turn length, and lexical diversity. Our experiments with FLAN-T5, GPT-2 and LLaMA-2 show the potential of demographic information, user emotions and implicit feedback to improve task completion and the factual consistency of the generated responses. Our human evaluation shows the positive impact on user acceptance. For future work, we are planning to improve the quality of annotations in FEDI and to generalize our collection framework to other tasks and domains to increase the representativeness of FEDI. We are also planning to extend FEDI to be multimodal, including images that represent emotions, as well as audio signals.
@@ 9 Limitations
The training and validation dialogues in FEDI were synthetically generated using GPT-3.5-Turbo. Thus, there is a probability that some data is unfaithful, hallucinated, or even harmful Kumar et al. (2023); Zhang et al. (2023); Malaviya et al. (2023). In addition, some of these dialogues may seem artificial and unnatural due to potentially conflicting demographic information, e.g., language style contradicting with age or occupation, although we did not observe this in our analysis. In our experiments, we analyzed the toxicity of responses generated by the models using the Perspective API and found it to be negligible ($\leq 3\%$ for most cases). However, we acknowledge that the detector may not capture all the potential harmful content. The generated data may also contain positive stereotypes, i.e., seemingly-harmless words or patterns that are offensive to specific demographic groups, which are not marked by the detector Cheng et al. (2023).
Another limitation concerns the feedback scenarios represented in the feedback dialogues. Since they were generated artificially, it is possible that some user reactions appear to be unnatural, counterintuitive, and maybe not even addressing the underlying generation error. However, we did not observe these issues in our analysis.
@@ Appendix A Task Descriptions
FEDI includes dialogues from two basic services provided by post offices, customer support for parcel shipping and topping up a prepaid SIM card. In customer support for parcel shipping, the task is to help the user choose the right shipping box and delivery option for their needs (given the weight of the goods to be sent and the destination). Topping up a prepaid SIM card is less of an advisory service since customers usually know how much they want to recharge, their telephone number, and which telephone provider they are with. Table 6 lists the slots for each task.
@@
Express or standard
delivery.
@@
and 3-4 weeks for inter-
national deliveries.
Top Up Prepaid SIM Card
@@
Vodafone, POSTE Mobile, …
@@
10 euro, 20 euro, 30 euro
@@
Choice or Recharge Phone.
@@
Table 6: Slot values for parcel shipping and SIM card top-up
In modern post offices, service robots or other virtual agents are more commonly used to provide such services in a self-service manner. However, if something goes wrong, e.g., the shipping boxes are empty or the credit card was rejected, customers must have the option of requesting assistance from a human employee. In this case, the customer is asked to tell the agent the type of service they need assistance with. In turn, the agent creates a ticket for a human employee and returns the ticket number. We consider this as a kind of subtask to the other tasks. However, since the slots are the same, Table 6 lists this as Request Ticket.
Table 6 lists the annotations for each task. As intent annotations, we use the following values: parcel_choice, recharge_a_prepaid_phone and request_ticket.
@@
For receptionist services, we only include one task in FEDI: access control. Table 7 shows the slots for this task.
@@
It is an essential task in hotels, office buildings, or other facilities with restricted access. Visitors usually need to register at the reception desk before being allowed to enter. As of today, electronic access controls (EAC) are more common than reception desks, especially in the case of office buildings, and they are becoming increasingly intelligent. In our case, we focus on a scenario in which a visitor has an appointment with an employee in an office building. To access the building, the visitor needs to provide the EAC with information about the appointment, e.g., the name of the host, date and time, and the room number. The EAC can then decide to grant access or to call the host for confirming the visitor’s identity. If necessary, the EAC can also provide additional safety information, e.g., hygiene guidelines. The intent annotation for utterances related to this task is building_access.
@@
For customer service in the insurance domain, we focus on question answering in the context of pet, health or heritage insurance, as we ll as bank transactions and account conditions. As a source, we use the insurance policies from POSTE Italiane, which are also available in English language [^16] POSTE Italiane Insurance Policies, last accessed 13. September 2023.. Table 8 lists the slots.
@@
In the past, customers called their insurance agent or visited their local bank branch for all questions related such topics. Today, it is more common to talk to chatbots or other service agents first and only in exceptional cases to human employees. Overall, we extracted 100 question-document pairs, i.e., questions paired with a paragraph that contains the answer, for bank transactions and account conditions, 78 for health, 84 for heritage, and 39 for pet insurance, from the POSTE documents. The intent annotation for utterances related to this task is question_answering.
@@
In the prompts for dialogue generation (see Appendix C), we instruct GPT-3.5 to have a separate turn at the beginning and ending of a dialogue in which both roles greet each other by also considering the generated background story. Such utterances have the intent annotation greeting. However, we do not consider this as a separate task in the sense of this work and do not evaluate it separately.

### v2

@@ 6 Quality Control for FEDI
   +User Feedback
96.9 95.3 83.5 77.2 60.2 0.02 44.4 22.1 88.2
GPT-2 Feedback-Free GPT-2 88.3 81.6 78.7 69.6 28.1 0.02 34.9 10.4 87.1
   +User Emotions 84.1 83.8 75.4 67.3 26.7 0.02 35.1 10.4 87.1
   +Demographic Info. 80.2 80.2 69.3 57.5 26.3 0.02 34.6 10.4 87.1
   +User Emotions
   +Demographic Info.
85.1 84.8 71.6 66.7 29.2 0.02 36.0 11.4 87.3
Feedback    +Generation Error 92.4 91.7 84.3 79.3 30.9 0.02 29.2 8.0 86.2
   +User Feedback 98.9 96.5 83.0 80.3 32.3 0.02 30.0 8.3 86.3
@@
   +User Feedback
94.7 93.3 88.0 80.8 35.5 0.01 30.3 9.7 86.4
Llama 2 Feedback-Free Llama 2 85.9 81.2 37.6 39.2 28.3 0.02 29.3 7.1 86.1
   +User Emotions 89.3 85.3 40.2 41.3 18.7 0.01 36.3 14.9 85.4
   +Demographic Info. 85.6 82.5 37.1 40.1 21.3 0.02 33.8 4.5 86.5
   +User Emotions
   +Demographic Info.
86.7 87.9 41.4 39.6 20.6 0.03 28.8 5.6 81.3
Feedback    +Generation Error 93.1 95.7 54.8 59.6 29.1 0.01 24.1 7.9 77.4
   +User Feedback 94.9 93.2 63.5 70.1 27.1 0.02 24.5 6.9 78.8
@@
   +User Feedback
82.4 83.6 46.3 47.2 33.5 0.03 25.0 9.2 80.1
Llama 2 In-Context 10.6 12.4 8.6 5.6 13.1 0.02 11.3 3.7 81.4
Table 4: Results of our main experiments (averaged over three runs). The best-performing models are printed in bold. Differences from the baselines that are greater than ±$\pm 1.0$ are colored green and red.
@@ 7 Experiments and Results
We conduct experiments using three models of different architecture and pretraining approaches, including Flan-T5 Chung et al. (2022) (780M), GPT-2 Radford et al. (2019) (780M) and Llama 2 Touvron et al. (2023b) (7B, plain pretrained version) [^12] The model weights for Flan-T5 and GPT-2 are available in the Huggingface Model Hub (last accessed 04 January 2024). Access to the weights for Llama 2 must be requested from Meta AI (last acessed 04 January 2024).. We first finetune the pretrained models using the feedback-free dialogues (baselines) and include the gold user emotions, demographic information and documents as part of the input sequences. We use the gold annotations of these signals to avoid bias from external components, such as emotion classifiers or document retrievers. For Llama 2, we only finetune the LoRA Hu et al. (2022) weights. We also provide in-context results for this model. We use the best feedback-free models for experiments with the feedback dialogues, in which we include the generation error and the user feedback utterance in the input sequence. We provide additional details in the Appendix, including hyperparameters and data configuration for feedback training (F.1) and input sequences (F.2). We also provide results for experiments using Llama 3 Dubey et al. (2024) (F.5), which was published shortly after we had completed our main experiments, and continual learning from feedback data (H).
7.1 Evaluation Metrics
@@
For task completion, we use the Inform and Success Budzianowski et al. (2018) metrics and additionally measure the accuracy of the predicted intent and slot values. To measure the factual consistency of the generated responses in question answering, we use $\text{Q}^{2}$ Honovich et al. (2021). Since the generation errors in FEDI include social aspects (see Appendix B), we use Perspective API to measure their toxicity, and F1-Score, BLEU(-n) Papineni et al. (2002) and BertScore Zhang et al. (2020) [^13] For Inform and Success, we use the implementation from Nekvinda and Dušek (2021) as a reference. For $\text{Q}^{2}$, we use the reference implementation which is available in GitHub. Perspective API is a free-to-use service provided by Google and Jigsaw. We measure the F1-Score based on the overlapping tokens in target and prediction. For BLEU Papineni et al. (2002) and BertScore Zhang et al. (2020), we use the implementation from the HuggingFace evaluation library v0.4.1 and with $n=4$ for BLEU (last access to all resources on 04 January 2024). to measure their generation accuracy.
7.2 Results
Table 4 shows the results achieved in the test dialogues. The feedback-free experiments show that including user emotions has the most positive impact. It improves the generation accuracy and factual consistency for Flan-T5 Chung et al. (2022) and GPT-2 Radford et al. (2019) (here in combination with demographic information), and the generation accuracy and task completion for Llama 2 Touvron et al. (2023b). The feedback experiments show improved task completion and factual consistency ($Q^{2}$) across all models. Regarding toxicity, we did not observe any negative impact from including generation errors, except for some outliers in Flan-T5 and Llama 2 (see Appendix F.6).
On the Influence of Generation Errors and User Feedback
We assume that the generation errors and user feedback used in training served as negative examples, helping the models to learn to generate more accurate intents and slots and responses that better reflect the knowledge documents (see Appendix F.3 for examples). An analysis on dialogue type level (Appendix F.4) also shows an increased generation accuracy for Flan-T5 and GPT-2, but only for question answering. We assume this is due to the knowledge document, which as part of the input sequence regulates the influence of error and feedback information Xu et al. (2023a, b); Ung et al. (2022). The responses generated for the other tasks still fit the context but often deviate from the target sequences. For Flan-T5 and GPT-2, this is reflected in the F1-Score, which measures word overlapping and is more affected than BLEU Papineni et al. (2002) and BertScore Zhang et al. (2020).
Additional Insights Regarding Llama 2
For Llama 2, we found that the generated responses often suffer from hallucinations (especially in the feedback-free dialogues). The reduced intent and slot accuracy also suggest a tendency towards hallucination here. We did not observe this in the experiments with Llama 3 Dubey et al. (2024) (Appendix F.5). We also observe that the results of the finetuned Llama 2 models are significantly higher than those of the in-context experiment (we included the task descriptions along with examples in the instruction), emphasizing the importance of finetuning for task-oriented and knowledge-grounded dialogues Zhang et al. (2024, 2023a).
@@ 7.2 Results
To investigate how humans perceive the impact of feedback training, we recruited 42 participants from Prolific [^14] Prolific is a widely used crowdsourcing platform for scientific research (last accessed 08 May 2024).. We asked them to rate the human likeness (Hum.), relevancy (Rel.), sociality (Soc.), engagement (Eng.), and factual consistency (Fact.) of the responses generated for 300 randomly sampled test dialogues in the feedback and feedback-free experiments highlighted in Table 4 (50 test dialogues from each experiment). We used a Likert scale from one to five for each attribute (with one as the lowest value). We received 40 valid submissions (we checked them manually in detail). Thus, each dialogue was rated by at least five participants. Table 16 shows the results. We provide more details on our rating scheme, annotator background and procedure in Appendix G.
Model
Hum.
Rel. Soc. Eng.
Fact.
IAA
@@
Feedb.-Free 3.41 4.12 4.66 3.56 4.12 0.25 0.15
Feedback 3.27 3.99 4.56 3.57 4.02 0.20 0.16
@@
Feedb.-Free 3.25 3.97 4.70 3.60 3.63 0.18 0.05
Feedback 4.02 3.88 4.58 3.52 3.64 0.21 0.11
Llama 2
Feedb.-Free 3.0 3.31 4.49 3.16 2.74 0.25 0.10
Feedback 3.12 3.87 4.64 3.54 3.69 0.23 0.10
Table 5: Results of our human evaluation. If statistically significant, they are printed in bold. (independent two-sample t-test, $p\leq 0.05$). We calculate IAA across all metrics (standard deviation in subscript) using Krippendorff’s Alpha Krippendorff (2006) [^16] We used SciPy v1.13.0 for the t-test (last accessed 08 April 2024). For Krippendorff’s Alpha, we used K-Alpha Calculator Marzi et al. (2024) (interval weighting)..
For Flan-T5 Chung et al. (2022) and GPT-2 Radford et al. (2019), annotators reported that the responses generated by the feedback models are more informative (which is not captured by the scores), but do not always cover the knowledge document as well as the responses from the feedback-free models (Flan-T5). They also reported them to be more direct and contain more counter-questions (which is actually desirable). This is often perceived as unfriendly, inattentive or disruptive and reflected in the slightly lower scores for relevancy and sociality (see Appendix G.2 for examples). For Llama 2 Touvron et al. (2023b), annotators reported some responses of the feedback-free model as illogical, unrelated to the dialogue context and factually incorrect. The responses generated by the feedback model were rated much higher, especially their relevancy and factual consistency. The IAA is rather low for most measures, which we attribute to their subjectivity and the diversity of annotators.
@@ 8 Conclusion
We introduce FEDI, the first English task-oriented and document-grounded dialogue dataset annotated with implicit user feedback, user emotions and demographic information. Our analysis shows the usefulness of our framework for generating feedback-annotated dialogues from various domains, and that FEDI is comparable to other related datasets. Our experiments show that learning from implicit user feedback improves task completion and factual consistency. Humans perceive the responses generated by feedback models as more informative (Flan-T5 and GPT-2), more relevant and more factually consistent (Llama 2). However, our results also show room for improvements in future work, e.g., the varying impact of learning from errors and feedback data on the generated responses and how they are perceived by humans.
@@ 9 Limitations
Taxonomies Used
The taxonomies used for generating implicit user feedback, user emotions and demographic information only reflect subsets of possible values. They are not exhaustive. For example, we do not consider educational background for demographic information, or other emotions than those that seemed meaningful to us in the context of this work. Our taxonomy of user emotions may differ from the original works.
Synthetically Generated Data
The training and validation dialogues in FEDI were generated using GPT-3.5. They are the result of a scripted generation procedure, and there is a probability that some data is unfaithful, hallucinated, or even harmful Kumar et al. (2023); Zhang et al. (2023b); Malaviya et al. (2023). Model-specific bias could also be a factor, which we haven’t investigated further. Although our analysis shows that the generated annotations are of high quality and we have invested a lot of effort in developing the instructions used, some values may be incorrect or inappropriate in the context, e.g., in the case of user emotions. This also applies to the slot and intent annotations, where analysis has shown that human annotators can react more flexibly to the dialogue background. In contrast, GPT-3.5 focuses entirely on the instruction and tends to return placeholder values in case of doubt. In addition, some of these dialogues may seem artificial and unnatural due to potentially conflicting demographic information, e.g., language style contradicting age or occupation. The same applies to the feedback scenarios represented in the feedback dialogues. Some user feedback may appear unnatural and counterintuitive and may not even relate to the underlying generation error. Although we conducted a fairly extensive human curation study in which we did not observe these issues, a more thorough review of the whole dataset would be required for a final assessment.
To solve feedback scenarios, we experimented with different ideas to incorporate the feedback into regenerating the affected system utterance. However, this led to unnatural and inconsistent dialogues, so we decided to use the naive approach described in the paper. As a result, the regenerated system utterances may not always directly reflect the feedback.
Toxicity Through Learning From Generation Errors
In our feedback experiments, we also use generation errors for learning. Since they also include social aspects, such as disrespectful or toxic response behavior, we used Perspective API to analyze the toxicity in generated responses. Although conspicuous responses were very rare, we acknowledge that the detector may not capture all the potentially harmful content. The generated data may also contain positive stereotypes, i.e., seemingly harmless words or patterns offensive to specific demographic groups, which are not marked by the detector Cheng et al. (2023).
Human Evaluation
We conducted the human evaluation as a crowdsourcing study and recruited 42 participants so that each dialogue was evaluated seven times. Some participants submitted their assessment far below the time limit, which is why we carefully checked each individual submission. Due to deviations from our rating scheme, we had to discard two submissions, which is why 100 of the 300 dialogues considered received fewer than seven ratings. Another limitation is the study design. We only considered the quality of the generated responses and not that of the generated slot and intent values. During the study, we found that our rating scheme has limitations as well. For example, hallucinations were not considered as a separate measure. Some annotators reported them as comments to the affected dialogues. However, the number was very small and we did not notice any additional cases when checking the submissions.
Acknowledgments
This work has been funded by the European Union under the Horizon Europe grant № 200009-57100412 (SERMAS).
@@ Appendix A Task Descriptions
FEDI includes dialogues from two basic services provided in post offices, customer support for parcel shipping and topping up a prepaid SIM card. In customer support for parcel shipping, the task is to help the user choose the right shipping box and delivery option for their needs (given the weight of the goods to be sent and the destination). Topping up a prepaid SIM card is less of an advisory service since customers usually know how much they want to recharge, their telephone number, and which telephone provider they are with. Table 6 lists the slots for each task.
@@
Express or standard delivery.
@@
and 3-4 weeks for international
deliveries.
Top Up SIM Card
@@
Vodafone, POSTE Mobile, … .
@@
10 euro, 20 euro, 30 euro.
@@
shipping or topping up a
prepaid SIM card.
@@
Table 6: Slot values for parcel shipping and topping up a prepaid SIM card.
In modern post offices, service robots or other virtual agents are more commonly used to provide such services in a self-service manner. However, if something goes wrong, e.g., the shipping boxes are empty or the credit card was rejected, customers must have the option of requesting assistance from a human employee. In this case, the customer is asked to tell the agent the type of service they need assistance with. In turn, the agent creates a ticket for a human employee and returns the ticket number. We consider this as a kind of subtask to the other tasks (Request Ticket in Table 6) and do not evaluate it separately.
@@
For receptionist services, FEDI only includes one task: access control. Table 7 shows the slots for this task.
@@
It is an essential task in hotels, office buildings, or other facilities with restricted access. Visitors usually need to register at the reception desk before being allowed to enter. As of today, electronic access controls (EAC) are more common than reception desks, especially in the case of office buildings, and they are becoming increasingly intelligent. In our case, we focus on a scenario in which a visitor has an appointment with an employee in an office building. To access the building, the visitor needs to provide the EAC with information about the appointment, e.g., the name of the host, date and time, and the room number. The EAC can then decide to grant access or to call the host for confirming the visitor’s identity. If necessary, the EAC can also provide additional safety information, e.g., hygiene guidelines.
@@
For customer service in the insurance domain, we focus on question answering in the context of pet, health and heritage insurance, as well as bank transactions and account conditions. As a source, we use the insurance policies from POSTE Italiane, which are also available in English language [^17] POSTE Italiane Insurance Policies, last accessed 13 January 2024.. Table 8 lists the slots.
@@
In the past, customers called their insurance agent or visited their local bank branch for all questions related to such topics. Today, it is more common to talk to chatbots or other service agents first and only in exceptional cases to human employees. Overall, we extracted 313 question-document pairs, i.e., questions paired with a paragraph that contains the answer, 19 for bank transactions, 93 for account conditions, 78 for health, 84 for heritage, and 39 for pet insurance, from the POSTE documents.
@@
In the prompts for dialogue generation (see Appendix C), we instruct GPT-3.5 to have a separate turn at the beginning and ending of a dialogue in which both roles greet each other by also considering the generated background story. However, we do not consider this as a separate task in the sense of this work and do not evaluate it separately.